when the price of a jar of jelly decreased from $4 to $2, suppliers of jelly decreased production from 250,000 jars to 150,000 jars. What is the price elasticity of supply in this case? What is the interpretation of the estimated elasticity?
1
Expert's answer
2020-04-06T09:53:06-0400

(150-250)/((150+250)/2)÷(2-4)/((2+4)/2)=-1/2÷(-2/3)=3/4=0.75 The supplu is not price elastic